A certain substance has a heat of vaporization of 31.17 kJ/mol. At what Kelvin temperature will the vapor pressure be 7.00 times higher than it was at 343 K?
Answer #1

T1 = 343 K

P2/P1 = 7

Hvap = 31.17 KJ/mol

= 31170 J/mol

we have below equation to be used:

ln(P2/P1) = (Hvap/R)*(1/T1 - 1/T2)

ln(7) = (31170.0/8.314)*(1/343.0 - 1/T2)

1.9459 = 3749.0979*(1/343.0 - 1/T2)

T2 = 417 K

Answer: 417 K